candyplays321:

Evaluate the expression 43 ÷ (7 − 3) × 2.

Joe348:

solve with PEMDAS

Joe348:

(7-3)=4

Alexis1415:

So we would start by using PEMDAS (Parentheses, Exponents, Multiplication, Division, Addition and Subtraction) 43/(4) times 2 10.75 x 2 = 21.5 or 86/8

Joe348:

\[43\div 4\times 2\]

Joe348:

@joe348 wrote:
\[43\div 4\times 2\]
\[21.5\]

candyplays321:

I am the stupidest person alive bro I have no clue what any of that means

Joe348:

P-Parenthesis E-Exponents M-Multiply D-Division A-Addition S-Subtraction

Joe348:

wym?

candyplays321:

nvm

Joe348:

well anyway you solve the equation in order of PEMDAS